# Learning Noise-Robust Stable Koopman Operator for Control with Hankel DMD

### v5

@@ V-E Overall Impact of Optimized Training Strategy
Finally, we compare the overall performance of our optimized configuration with the baseline configuration, as shown in 14 and 15. 14 shows the prediction performance comparison for the Van der Pol system and 15 shows the model predictive costs comparison for the Cartpole system.
For the Van der Pol system 14, the baseline setup utilized the Standard Koopman Operator approach with a fixed maximum roll-out duration of 99, without any learning rate scheduler and maintained a constant learning rate of 0.01. In contrast, our optimized configuration employed a dissipative Koopman operator alongside a progressive roll-out strategy, starting with a roll-out length of 1 and geometrically increasing by a factor of 2 every 200 epochs until reaching the maximum roll-out length. Additionally, we implemented a triangular mode cyclic learning rate scheduler, with the maximum learning rate set to 0.1, halving every 500 epochs, and a base learning rate of 0.01. Both configurations used the Adam optimizer. In general, the optimized configuration resulted in a 39% improvement in training error and a 37.6% improvement in testing error compared to baseline. This significant reduction in error highlights how the combined components of the optimized configuration effectively enhance the model’s performance.
In the case of the Cartpole system (15, the baseline strategy used the Adam optimizer with a constant roll-out duration of 16 and a constant learning rate. The Standard Koopman Operator approach was used throughout the training. Conversely, the optimized strategy employed the Adam optimizer for the initial 1500 epochs, followed by fine-tuning using the LBFGS optimizer. The duration of the roll-out was progressively increased from 1 to 16, a cyclic learning rate was implemented, and a dissipative Koopman operator was used. The optimized training strategy outperforms the baseline training strategy in different initial conditions in multiple iterative training data augmentations.
@@
Figure 14: Comparison of training and testing errors between the optimized and baseline configurations across various noise levels. The optimized configuration shows a 39% improvement in training error and a 37.6% improvement in testing error.
Figure 15: Comparison of Model predictive costs for Cartpole problem for 4 different initial conditions between the optimized and baseline configurations.
@@ Acknowledgement
The authors thank the Center for Computational Innovations (CCI) for providing computational resources during the early stages of this research. Numerical experiments are performed using the computational resources granted by NSF-ACCESS for the project PHY240112 and that of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility using NERSC award NERSC DDR-ERCAP0030714.

### v6

@@ V-E Overall Impact of Optimized Training Strategy
To assess the combined effect of the above-mentioned training strategies, we compare the overall performance of our optimized configuration with ”baseline” configurations for Van der Pol system fig. 14 and Cartpole system fig. 15. Figure 14 shows the prediction performance comparison for the Van der Pol system and fig. 15 shows the model predictive costs comparison for the Cartpole system over iterative data augmentation.
For the Van der Pol system fig. 14, the baseline configuration utilized the Standard Koopman Operator approach with a fixed maximum roll-out duration of 99, without any learning rate scheduler and maintained a constant learning rate of 0.01. In contrast, our optimized configuration employed a dissipative Koopman operator alongside a progressive roll-out strategy, starting with a roll-out length of 1 and geometrically increasing by a factor of 2 every 200 epochs until reaching the maximum roll-out length. Additionally, we implemented a triangular mode cyclic learning rate scheduler, with the maximum learning rate set to 0.1, halving every 500 epochs, and a base learning rate of 0.01. Both configurations used the Adam optimizer. In general, the optimized configuration resulted in a 39% improvement in training error and a 37.6% improvement in testing error compared to baseline. This significant reduction in error highlights how the combined components of the optimized configuration effectively enhance the model’s performance.
In the case of the Cartpole system (fig. 15, the baseline configuration used the Adam optimizer with a constant roll-out duration of 16 and a constant learning rate. The Standard Koopman Operator approach was used throughout the training. Conversely, the optimized strategy employed the Adam optimizer for the initial 1500 epochs, followed by fine-tuning using the LBFGS optimizer. The duration of the roll-out was progressively increased from 1 to 16, a cyclic learning rate was implemented, and a dissipative Koopman operator was used. The optimized training strategy outperforms the baseline training strategy in different initial conditions in multiple iterative training data augmentations.
@@
Figure 14: Comparison of model prediction performance for the Van der Pol system. The “baseline” refers to the standard Koopman operator trained with fixed rollout length (R=99), no learning rate scheduling, and constant learning rate of 0.01. The “optimized” strategy incorporates a dissipative Koopman operator, progressive rollout, and cyclic learning rate schedule. The optimized configuration shows a 39% improvement in training error and a 37.6% improvement in testing error.
Figure 15: Model predictive control performance for the Cartpole system. The “baseline” configuration uses the standard Koopman operator with fixed rollout length (R=16), constant learning rate, and no learning rate scheduling. The “optimized” configuration employs a dissipative Koopman operator, progressively increasing rollout length, a cyclic learning rate schedule, and fine-tuning with the LBFGS optimizer after initial training with Adam. Performance is evaluated across multiple iterative training data augmentations and initial conditions.
@@ Acknowledgement
This work used NPL at RPI through allocation AGRM from the Center for Computational Innovations during the early stages of this research. Numerical experiments were performed on the FASTER and Anvil supercomputers at TAMU and Purdue, respectively, using computational resources granted by the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program under allocation PHY240112 (supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296) and by the National Energy Research Scientific Computing Center (NERSC), a DOE Office of Science User Facility, under award DDR-ERCAP0030714. The authors appreciate the use of the research computing resources of the Empire AI Consortium, Inc. with support from Empire State Development of the State of New York, the Simons Foundation, and the Secunda Family Foundation.